40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Message ID: 1120     Entry time: Thu Mar 14 22:39:49 2013
Author: Evan 
Type: DailyProgress 
Category: RefCav 
Subject: South cavity locked on TEM00 

 [Tara, Evan]

Tara and I spent today and Tuesday laying down the new optical path for the south cavity according to the layout in entry 1100. After quite a bit of periscope steering (mostly on Tara's part), we got flashes of TEM00 and subsequently were able to lock the laser's fast frequency control servo to the cavity resonance. Getting near the resonance requires a bit of dialing around with the set point on laser temperature servo; the values that worked today were 512 on the coarse adjust and 894 on the fine adjust.

From here, we should

  1. make a new mount for the post-PMC Faraday isolator, since the existing mount is about 32 mils too high (and hence the FI has not yet been placed in the beam path);
  2. fine-tune the steering into the cavity;
  3. optimize the mode-matching into the cavity;
  4. measure the cavity finesse;
  5. pick a different lens for matching into the RFPD, since the current RFPD path length is too long; and
  6. fine-tune the demodulation phase for the loop.
ELOG V3.1.3-